# Hole-spin qubits in germanium beyond the single-particle regime

### v1

@@ Abstract
The intense simulation activity on hole-spin qubits in Ge has focused so far on singly occupied quantum dots. Here, we theoretically investigate three-hole qubits in germanium and show that their performances compare well with those of their single-hole counterparts, both in strained and unstrained systems. In particular, enhancements of the gating times by two orders of magnitude are obtained in slightly elongated quantum dots. We show that, depending on the confinement anisotropy, such an enhancement is driven either by the occupation of excited spin-orbitals (Pauli principle), or by a Coulomb-induced charge rearrangement.
Figure 1: (a) Schematic representation of a single-hole state (left) and a three-hole state (right). (b) Schematic representations of the device that we consider in this work, which will either represent an unstrained MOS device, or a strained Ge/Si1-xGex heterostructure. In both cases, the Ge channel has a width $L$ along the $z$ direction. A potential is applied through a metallic top gate, forming a QD that confines the holes. The profile $V_{\perp}$ of the confinement along the $z$ direction is sketched in the rightmost part of panel (b).
@@ Introduction.
The vast majority of the literature on hole-spin qubits focuses on single-particle encodings. Multiple holes, localized in singly-occupied QDs, have also been considered for alternative qubit encodings [10, 59, 60, 61, 62, 63, 20, 64]. On the other hand, an odd number of holes in a single QD enables qubit encodings and manipulation schemes analogous to those implemented in the single-particle case [65, 66, 67]. Such a multi-hole qubit encoding has a twofold interest. First, from an experimental point of view, it relaxes the stringent requirement of achieving the single-occupation regime. Second, and most importantly, it might result in larger and more local couplings to the control gates and, ultimately, in improved figures of merit. However, despite their past and recent occurrence in experiments [11, 12, 44, 45, 68, 69], few-particle effects in semiconductor hole-spin qubits remain relatively unexplored at the theoretical level.
@@
In this Letter, we provide a theoretical study of multi-hole qubits in Ge QDs, implemented in unstrained MOS-like devices and in Ge///Si1-xGex heterostructures (Fig. 1). Focusing on the case of three holes, we show that the main figures of merit are always at least comparable to those obtained for single holes, and significantly improved in specific regimes. More specifically, the Rabi frequency of the three-hole qubit (THQ) is shown to exceed that of a single-hole qubit (SHQ) by up to two orders of magnitude in slightly squeezed QDs. This occurs without a significant modification of the Larmor frequency or loss of the qubit coherence. As the comparison of the fully interacting THQ with its noninteracting counterpart shows, this gain is largely of single-particle origin, as it mostly results from the occupation of the excited orbitals dictated by the Pauli principle. However, a Coulomb-driven charge rearrangement manifests itself in a significant peak in the Rabi frequencies of the interacting THQ. On the other hand, the strain, while changing the absolute values of the manipulation frequencies, does not substantially affect the relative performances of the single- and multi-hole qubits.
Figure 2: Comparison between (a-b) the Larmor and (c-f) the Rabi frequencies of the SHQ (left panels) and THQ (right), with fixed $B=0.05$ T and $\left|\delta\bm{E}_{\rm R}\right|=1$ mV///nm. The frequencies are plotted as a function of $\hbar\omega_{x}$, and for an oscillating electric field oriented along the $x$ (c, d) or $y$ (e, f) directions. The curves correspond to $\theta=m\times 10\degree$, where $m$ ranges from 00 (red) to 9 (black).
Qubit model.
@@
Single-hole states are derived through the numerical diagonalization of the 6-band $\bm{k}\cdot\bm{p}$ Hamiltonian resulting from the Luttinger-Kohn (LK) envelope-function approach [70, 71, 53, 72]. The 6 bands needed to describe single-hole states in Ge close to the top of the valence band form a $J=3/2$ quadruplet and a $J=1/2$ doublet, with $J$ being the total atomic angular momentum associated with the Bloch states. These two multiplets are split at the $\Gamma$ point by SOC, the gap being $\Delta_{\rm SO}=290$ meV. The $J=3/2$ quadruplet, degenerate at $\Gamma$, is split into the HH ($|M|=3/2$) and the LH ($|M|=1/2$) doublets at finite $k$. The band structure around the $\Gamma$ point is determined by the LK parameters, which, for Ge, are: $\gamma_{1}=13.38$, $\gamma_{2}=4.24$ and $\gamma_{3}=5.69$. The reference frame, defined by the unit vectors $\bm{u}_{x}$, $\bm{u}_{y}$, and $\bm{u}_{z}$, is defined with respect to the crystallographic axes as follows: $\bm{u}_{x}\parallel[110]$, $\bm{u}_{y}\parallel[\bar{1}10]$, $\bm{u}_{z}\parallel[001]$.
@@
The confining potential varies smoothly with respect to the unit-cell scale, and has the form
@@
Here, $m_{0}$ is the free-electron mass, while $\hbar\omega_{y}$ (set to 6 meV) and $\hbar\omega_{x}$ (from 3 to 5.5 meV) are the characteristic energy scales associated with the anisotropic harmonic confinement in the $xy$ plane. The out-of-plane confinement defines a well with sharp barriers, accounting for the band offset between Ge and the barrier material:
@@
Here, $L=10\,$nm is the well width, $W$ is the barrier height, and $s=0.02\,$nm a smoothing factor. The dc electric field $e_{\perp}=1\,$meV///nm, produced by the top gate voltage, is attractive for holes.
The coupling to the magnetic field is introduced into the single-hole Hamiltonian through the Zeeman-Bloch term and the Peierls substitution [52, 35], with the electromagnetic vector potential expressed in the Coulomb gauge. The magnetic field is given in polar form, $\bm{B}\equiv B\,(\sin\theta\,\cos\phi,\,\sin\theta\,\sin\phi,\,\cos\theta)$, with $B=0.05\,$T, fixed in-plane angle $\phi=45\degree$, and variable out-of-plane angle $\theta$.
Leveraging the high excitation energies along the $z$ direction and in order to provide an accurate description of in-plane correlations at the few-particle level, we partially freeze the motion in the vertical direction and identify — for each band — the $z$ component of the wave function with the ground state of the respective $z$-dependent Hamiltonian term. The few-particle (three-hole) states are computed by diagonalizing the interacting Hamiltonian, within the full configuration interaction approach. Here, the Coulomb interaction is included within the intraband approximation [73], and reduced to an analytical 2D potential, consistently with the character of the single-particle states. The three-hole Slater determinants are built from a set of 64 single-particle states, which ensures the convergence of the Larmor and Rabi frequencies (further details on the method are provided in the Supplemental Material (SM) [74]).
The magnetic field breaks the Kramers degeneracy within the ground doublet, and induces the energy splitting $\Delta^{(n)}_{10}$ between the states $\left|0^{(n)}\right\rangle$ and $\left|1^{(n)}\right\rangle$, which define the $n$-hole qubit; here we compare $n=1$ and $n=3$. This splitting can equivalently be expressed in terms of the Larmor frequency $f^{(n)}_{\rm L}\equiv\Delta^{(n)}_{10}/h\equiv g^{(n)}\mu_{\rm B}B/h$, where $g^{(n)}(\theta,\phi)$ is the effective $g$-factor, which depends on the orientation of $\bm{B}$. In the following, we specifically refer to $g^{(n)}_{\perp}$ and $g^{(n)}_{\parallel}$, corresponding to out-of-plane ($\theta=0^{\circ}$) and in-plane ($\theta=90^{\circ}$) directions, respectively.
$\hbar\omega_{x}=3\,$meV (elongated) $\theta=0^{\circ}$ (out-of-plane) $\theta=90^{\circ}$ (in-plane)
$g_{\perp}$ $f^{(n)}_{{\rm R},x}$ (MHz) $f^{(n)}_{{\rm R},y}$ (MHz) $g_{\parallel}$ $f^{(n)}_{{\rm R},x}$ (MHz) $f^{(n)}_{{\rm R},y}$ (MHz)
SHQ 17.9 203 99 0.13 1.3 0.32
THQ (non interacting) 15.9 453 48 0.28 1.1 0.84
THQ (interacting) 17.1 116 17 0.18 0.51 0.11
$\hbar\omega_{x}=5.5\,$meV (quasi-circular) $\theta=0^{\circ}$ (out-of-plane) $\theta=90^{\circ}$ (in-plane)
$g_{\perp}$ $f^{(n)}_{{\rm R},x}$ (MHz) $f^{(n)}_{{\rm R},y}$ (MHz) $g_{\parallel}$ $f^{(n)}_{{\rm R},x}$ (MHz) $f^{(n)}_{{\rm R},y}$ (MHz)
SHQ 17.3 18 40 0.16 0.12 0.25
THQ (non interacting) 9.4 300 160 0.54 2.1 9.5
THQ (interacting) 14.1 228 128 0.33 0.19 2.5
Table 1: Values of the $g$ factors and of the Rabi frequencies of the unstrained single- and three-hole qubits (interacting and non-interacting cases). The quantities are computed for $\hbar\omega_{y}=6\,$meV, $B=0.05$ T, and $\left|\delta\bm{E}_{\rm R}\right|=1$ mV///nm.
@@
$f_{\rm R}^{(n)}=\frac{e}{h}\Big{|}\delta\bm{E}_{\rm R}\cdot\langle 0^{(n)}|\hat{\bm{r}}|1^{(n)}\rangle\Big{|}\,,$ (3)
@@
where $\hat{\bm{r}}$ is the hole position operator, and $|\delta\bm{E}_{\rm R}|=1\,$mV/nm is the amplitude of the spatially homogeneous oscillating electric field. Hereafter, we specifically discuss the frequencies $f_{{\rm R},\alpha}^{(n)}$, corresponding to an electric field oriented along the coordinate axis $\alpha$.
Three- vs single-hole spin qubits.
The comparison between single- and multi-hole encodings is mainly defined in terms of the respective Larmor and Rabi frequencies. Concerning the former ones, no significant difference is observed between SHQs and THQs [Fig. 2(a, b)]. Both $f_{\rm L}^{(1)}$ and $f_{\rm L}^{(3)}$ are characterized by a weak dependence on the in-plane aspect ratio of the dot (i.e. on $\omega_{x}$, for a fixed $\omega_{y}$), and by a dependence on $\theta$ that, for $|\theta-90^{\circ}|\gg 1^{\circ}$, reads: $f^{(n)}_{\rm L}(\theta)\approx f^{(n)}_{\rm L}(0)\,\cos\theta$. This relation results from the strong $g$ factor anisotropy [5], which is only weakly attenuated in the few-particle case (see Table 1).
While the Rabi frequencies of the THQ and of the SHQ have a similar dependence on the orientation of $\bm{B}$, they differ remarkably in their dependence on the spatial confinement and specifically on the in-plane aspect ratio [Fig. 2(c-f)]. In particular, in the quasi-circular-QD regime ($\omega_{x}\approx\omega_{y}$), the THQ performs largely better than the SHQ for all orientations of $\bm{B}$, whereas the SHQ Rabi frequencies are higher than those of its three-hole counterpart in the elongated-dot regime ($\omega_{x}\ll\omega_{y}$). The differences between the two cases become even more significant as the orientation of $\delta\bm{E}_{\rm R}$ changes from the weak (${\bm{u}}_{x}$) to the strong-confinement direction (${\bm{u}}_{y}$).
A striking feature of Fig. 2(d, f) is the development of a peak in the graphs of the THQ Rabi frequencies at $\hbar\omega_{x}\approx 3.8$ meV. This is a genuinely many-body effect that corresponds to a transition between different classical equilibrium configurations [75] of the quasi-2D 3-hole system. In fact, the classical potential energy is minimized by a linear arrangement of the particles along the $x$ axis for $\omega_{x}<\omega_{x,{\rm th}}\equiv\omega_{y}\sqrt{5/12}$, and by a triangular arrangement for $\omega_{x}>\omega_{x,{\rm th}}$. For $\hbar\omega_{y}=6$ meV, one has that $\hbar\omega_{x,{\rm th}}\approx 3.87$ meV, in good agreement with the position of the peaks (more details can be found in the SM [74]).
Role of Coulomb interaction.
@@
The question arises as to whether the differences between the THQ and the SHQ are mostly due to Coulomb interaction and to the resulting charge localization, or to the occupation of the excited single-particle states, imposed by the Pauli principle. In order to discriminate between these two cases, we compare the values of the $g$ factors and of the Rabi frequencies obtained for the fully-interacting and for the noninteracting regimes (Table 1).
Both in the quasi-circular and in the elongated QDs, the Coulomb interaction reduces the $g$-factor anisotropy (albeit slightly) and decreases the differences between the $g$ factors in the single- and the three-qubit cases. This points to the fact that such differences are dominated by Pauli-principle effects [76], except for the region where the system undergoes a Coulomb-driven charge redistribution [see the peaks in Fig. 2(d, f)].
The Coulomb interaction has a significant impact on the Rabi frequencies. In moving from the noninteracting to the fully interacting regimes, all of the $f_{{\rm R},\alpha}^{(3)}$ decrease. In fact, for some of the considered field orientations, the noninteracting THQ displays Rabi frequencies that are larger than those of the SHQ even in the elongated-QD case. This clearly indicates that the advantage of the THQ with respect to the SHQ is of single-particle origin (i.e. is related to the occupation of excited orbitals), and tends to be reduced by the Coulomb interaction.
$\hbar\omega_{x}=3\,$meV (elongated) $\theta=0^{\circ}$ (out-of-plane) $\theta=90^{\circ}$ (in-plane)
$f^{(n)}_{{\rm R},x}$ $f^{(n)}_{{\rm R},y}$ $f^{(n)}_{{\rm R},x}$ $f^{(n)}_{{\rm R},y}$
SHQ 116 55 0.48 0.053
THQ 68 8.6 0.16 0.099
$\hbar\omega_{x}=5.5\,$meV (quasi-circular) $\theta=0^{\circ}$ (out-of-plane) $\theta=90^{\circ}$ (in-plane)
$f^{(n)}_{{\rm R},x}$ $f^{(n)}_{{\rm R},y}$ $f^{(n)}_{{\rm R},x}$ $f^{(n)}_{{\rm R},y}$
SHQ 10 22 0.036 0.07
THQ 140 72 0.017 0.85
Table 2: Values of the Rabi frequencies (expressed in MHz) of the strained single- and three-hole qubits. The quantities are computed for $\hbar\omega_{y}=6\,$meV, $B=0.05$ T, and $\left|\delta\bm{E}_{\rm R}\right|=1$ mV///nm.

### v2

# arXiv:2505.02449v2[cond-mat.mes-hall] 01 Mar 2026

@@ Abstract
The intense simulation efforts on hole-spin qubits in germanium have so far focused primarily on singly occupied quantum dots. Here, we theoretically investigate three-hole qubits in germanium and demonstrate that their performance can rival that of single-hole qubits in both strained and unstrained systems. In particular, we find that — in the widely used quasi-circular geometry — a three-hole qubit encoding can yield enhancements of the Rabi frequencies of up to two orders of magnitude and a large advantage also in terms of quality factors.
Figure 1: (a) Pictorial representation of a single-hole state (left) and a three-hole state (right) in a QD elongated along the $x$ direction. (b) Schematic view of a MOS device, where the Ge channel has a width $L$ along the $z$ direction. The bias applied to the metal top gate determines the in-plane confining potential $V_{\parallel}$ and the linear (in $z$) contribution in the out-of-plane confining potential $V_{\perp}$.
Figure 2: In-plane particle distribution $N_{z}(x,y)$ for the ground state $|\Phi_{0}\rangle$ of the three-hole system, obtained by integrating the particle density $n({\bf r})$ along the $z$ direction. The four panels display: (a) the function $N_{z}(x,y)$; (b,c) the heavy-hole contributions; (d) the sum of the two light-hole contributions. All the plotted functions are normalized to 1, the isolines correspond to values that are multiples of 0.1. The values of the parameters are: $\hbar\omega_{x}=5\,$meV, $\hbar\omega_{y}=6\,$meV, B=0.05B=0.05\,T and $\theta=0$ (parallel field orientation).
@@ Introduction.
The vast majority of the literature on hole-spin qubits focuses on single-particle encodings. Multiple holes, localized in singly-occupied QDs, have also been considered for alternative qubit encodings [10, 59, 60, 61, 62, 63, 20, 64]. On the other hand, an odd number of holes in a single QD enables qubit encodings and manipulation schemes analogous to those implemented in the single-particle case [65, 66, 67]. Such a multi-hole single-dot qubit has a twofold interest. First, from an experimental point of view, it relaxes the stringent requirement of achieving the single-occupation regime. Second, and most importantly, it might result in larger and more local couplings to the control gates and, ultimately, in improved figures of merit [68]. However, despite their past and recent occurrence in experiments [11, 12, 44, 45, 69, 70], few-particle effects in semiconductor hole-spin qubits remain relatively unexplored at the theoretical level.
@@
In this Letter, we provide a theoretical study of multi-hole qubits in Ge QDs, implemented in unstrained MOS-like devices and in Ge//Si1-xGex heterostructures (Fig. 1). Focusing on the case of three holes, we show that the main figures of merit are always at least comparable to those obtained for single holes, and significantly improved in specific regimes. More specifically, the Rabi frequency of the three-hole qubit (THQ) is shown to exceed that of a single-hole qubit (SHQ) by up to two orders of magnitude in quasi-circular QDs. This occurs without a significant modification of the $g$ factor or loss of the qubit coherence under charge noise. As the comparison of the fully interacting THQ with its noninteracting counterpart shows, this gain is caused, for the largest part, by the occupation of the excited orbitals dictated by the Pauli principle, the hole-hole interaction giving a quantitatively smaller (albeit far from negligible) contribution in that regime. Strain changes the absolute values of the manipulation frequencies, but does not substantially affect the relative performances of the single- and multi-hole qubits.
Quantum-dot model.
@@
Single-hole states are derived through the numerical diagonalization of the 6-band $\bm{k}\cdot\bm{p}$ Hamiltonian resulting from the Luttinger-Kohn (LK) envelope-function approach [71, 72, 53, 73]. The 6 bands needed to describe single-hole states in Ge close to the top of the valence band form a $J=3/2$ quadruplet and a $J=1/2$ doublet, with $J$ being the total atomic angular momentum associated with the Bloch states at the $\Gamma$ point. These two multiplets are split at $\Gamma$ by spin-orbit coupling (SOC), which induces a gap $\Delta_{\rm SO}=290$ meV. The $J=3/2$ quadruplet, degenerate at $\Gamma$, is split into the HH ($|M|=3/2$) and the LH ($|M|=1/2$) doublets at finite $k$. The band structure around the $\Gamma$ point is determined by the LK parameters, which, for Ge, are: $\gamma_{1}=13.38$, $\gamma_{2}=4.24$ and $\gamma_{3}=5.69$. The reference frame, specified by the unit vectors $\bm{u}_{x}$, $\bm{u}_{y}$, and $\bm{u}_{z}$, is defined with respect to the crystallographic axes as follows: $\bm{u}_{x}\parallel[110]$, $\bm{u}_{y}\parallel[\bar{1}10]$, $\bm{u}_{z}\parallel[001]$.
QD1 $g^{(1)}_{\theta}$ $g^{(3)}_{\theta}$ $f^{(3)}_{R,x}$ $f^{(3)}_{R,y}$
$\theta=0^{\circ}$ 8.67 7.99 (6.25) 46.8 (239) 9.24 (27.3)
$\theta=90^{\circ}$ 0.121 0.0848 (0.104) 7.29 (3.93) 2.07 (0.0775)
QD2 $g^{(1)}_{\theta}$ $g^{(3)}_{\theta}$ $f^{(3)}_{R,x}$ $f^{(3)}_{R,y}$
$\theta=0^{\circ}$ 8.00 4.88 (4.34) 394 (322) 313 (175)
$\theta=90^{\circ}$ 0.114 0.220 (0.321) 27.1 (23.9) 11.3 (2.04)
Table 1: Values of the $g$ factors and of the Rabi frequencies (expressed in MHz) for the unstrained single- and three-hole qubits, for $\hbar\omega_{y}=6\,$meV, $B=0.05$ T, and $\left|\delta\bm{E}_{\rm R}\right|=1$ mV//nm. The value of $\hbar\omega_{x}$ is 3 and 5.55.5\,meV for QD1 and QD2, respectively. In parentheses we report the Rabi frequencies of the noninteracting THQ.
@@
The confining potential that defines the dot model varies smoothly with respect to the unit-cell scale, and is specifically given by
@@
Here, $m_{0}$ is the free-electron mass, while $\hbar\omega_{y}$ (set to 6 meV) and $\hbar\omega_{x}$ (from 3 to 5.5 meV) are the characteristic energy scales associated with the anisotropic harmonic confinement in the $xy$ plane. The out-of-plane confinement defines a quantum well, resulting from the band offset between Ge and the barrier material:
@@
Here, L=10L=10\,nm is the well width, $W$ is the barrier height, and s=0.02s=0.02\,nm a smoothing factor. The dc electric field $e_{\perp}=1\,$meV//nm, possibly associated to a top gate voltage, is attractive for holes.
The coupling to the magnetic field is introduced into the single-hole Hamiltonian through the Zeeman-Bloch term and the Peierls substitution [52, 35], with the electromagnetic vector potential expressed in the Coulomb gauge. The magnetic field is given in polar form, $\bm{B}\equiv B\,(\sin\theta\,\cos\phi,\,\sin\theta\,\sin\phi,\,\cos\theta)$, with B=0.05B=0.05\,T, $\phi=45\degree$, and a variable polar angle $\theta$.
The few-particle (three-hole) eigenstates are computed by diagonalizing the interacting Hamiltonian, within the full configuration interaction approach [74, 75]. The three-hole Slater determinants are built from a set of 64 single-particle states, which ensures the convergence of the relevant quantities. Further details on the method are provided in the Supplemental Material (SM) [76].
Single- and three-hole spin qubits.
The magnetic field breaks the Kramers degeneracy within the ground doublet, and induces the energy splitting $\Delta^{(k)}_{10}$ between the states $\left|0^{(k)}\right\rangle$ and $\left|1^{(k)}\right\rangle$, which define the SHQ ($k=1$) and the THQ ($k=3)$. This splitting can equivalently be expressed in terms of the $g$ factor $g^{(k)}_{\theta}=\Delta^{(k)}_{10}/\mu_{\rm B}B$, which depends on the orientation of $\bm{B}$ with respect to the $z$ axis, defined by $\theta$.
Figure 3: Rabi frequencies $f^{(n)}_{{\rm R},\alpha}$ of the $n$-hole qubit as functions of $\hbar\omega_{x}$, with $n=1$ (red curves) and $n=3$ (blue symbols), when the oscillating electric field (of amplitude $\left|\delta\bm{E}_{\rm R}\right|=1$ mV//nm) is oriented along the $x$ (a, c) or $y$ direction (b, d). The magnetic field amplitude is $|\bm{B}|=0.05$ T; its direction is given by $\theta=0\degree$ (a, b) or $\theta=90\degree$ (c, d), with $\phi=45\degree$.
@@
$f_{\rm R}^{(k)}=\frac{e}{h}\Big|\delta\bm{E}_{\rm R}\cdot\langle 0^{(k)}|\hat{\bm{r}}|1^{(k)}\rangle\Big|\,$, (3)
@@
where $\hat{\bm{r}}$ is the hole position operator, and $|\delta\bm{E}_{\rm R}|=1\,$mV/nm is the amplitude of the spatially homogeneous oscillating electric field.
As a first means to characterize the three-hole ground state $|0^{(3)}\rangle$, we compute the particle density
$n(\bm{r})=\sum_{\beta}n_{\beta}(\bm{r})=\sum_{\beta}\langle 0^{(3)}|\hat{\psi}_{\beta}^{\dagger}(\bm{r})\,\hat{\psi}_{\beta}(\bm{r})|0^{(3)}\rangle\,$, (4)
where $\beta$ runs over the six hole bands and the field operator $\hat{\psi}_{\beta}(\bm{r})$ annihilates a hole belonging to the band $\beta$ at the position $\bm{r}$. Being the particle density along the vertical direction weakly correlated with that in the planar directions [76], we focus on the in-plane particle distribution, represented by $N_{z}(x,y)\equiv\int dz\,n(\bm{r})$ (Fig. 2). The three holes clearly display a linear arrangement along the weak-confinement direction ($x$), characterized by the presence of three peaks [panel (a)], which become increasingly resolved for decreasing values of $\omega_{x}$ [76]. Besides, the minority [$M=+3/2$, panel(b)] and majority [M=3/2M=-3/2, panel(c)] HH components of $N_{z}(x,y)\equiv\int dz\,n(\bm{r})$ are clearly related to the central and external peaks, respectively: the ground state of the THQ is thus characterized by an “antiferromagnetic" ordering in the pseudospin defined by the HH component ($M=\pm 3/2$). The in-plane distribution of the LH components [panel (d)], whose weight represent less than 1% of the total, display markedly different features.
[…]
QD1 $f^{(1)}_{R,x}$ $f^{(1)}_{R,y}$ $\tau^{(1)}$ $f^{(3)}_{R,x}$ $f^{(3)}_{R,y}$ $\tau^{(3)}$
$\theta=0^{\circ}$ 57.5 27.0 5.78 23.0 5.30 5.92
$\theta=90^{\circ}$ 0.276 0.0804 1480 0.0941 0.157 1130
QD2 $f^{(1)}_{R,x}$ $f^{(1)}_{R,y}$ $\tau^{(1)}$ $f^{(3)}_{R,x}$ $f^{(3)}_{R,y}$ $\tau^{(3)}$
$\theta=0^{\circ}$ 4.48 10.7 5.93 63.7 33.9 4.50
$\theta=90^{\circ}$ 0.0223 0.0494 752 0.301 0.349 3690
[…]
While the $g$ factors of the THQ and of the SHQ show a very similar dependence on the orientation of $\bm{B}$, the Rabi frequencies differ remarkably in the two cases (Fig. 3). In fact, for all the considered orientations of $\bm{B}$ and of the oscillating electric field $\delta\bm{E}_{\rm R}$, the Rabi frequency of the THQ displays a non-monotonic dependence on the dot aspect ratio ($\omega_{x}$). Moreover, the $f_{R,\alpha}^{(3)}$ are characterized by the presence of pronounced maxima, and by regions corresponding to a large enhancement with respect to the SHQ, mostly — but not exclusively — concentrated in the quasi-circular dot regime ($\hbar\omega_{x}\gtrsim 5\,$meV). Sharp peaks in the Rabi frequencies of the THQ show up at $\hbar\omega_{x}\approx 3.8\,$meV. These features correspond to, and are most likely related to, the presence of a narrow anticrossing between the first and second excited doublet [76].
@@
The question arises as to whether these striking differences between the THQ and the SHQ are mainly due to Coulomb interaction or to the occupation of excited single-particle states, imposed by the Pauli principle [77]. In order to discriminate between these two cases, we compare the above values of the Rabi frequencies with those obtained for the noninteracting three-hole system ($\bar{f}_{R,\alpha}$). The difference between the two cases is always remarkable and shows that the values of the Rabi frequencies are strongly affected by the Coulomb interactions. Whether the Rabi frequencies of the THQ are larger or smaller that those of the noninteracting three-hole system depends on the orientation of the magnetic field and on the degree of elongation of the QD (Table 1).
Figure 4: Dephasing time scale $\tau$ as a function of $\hbar\omega_{x}$ (with $\hbar\omega_{y}=6\,$meV), for a magnetic field of intensity B=0.05B=0.05\,T oriented (a) parallel ($\theta=0^{\circ}$) or (b) perpendicular ($\theta=90^{\circ}$) to the $z$ axis. The charge-noise induced electric field is oriented along the $z$ direction, with $\left|\bm{E}_{\rm cn}\right|=10^{-2}$ mV/nm.
Qubit dephasing.
Random fluctuations in the energy gap $\Delta_{10}$, induced by charge noise, are generally regarded as the main source of decoherence for hole-spin qubits [2]. The dephasing time scale $\tau$ can thus be defined as:
$\frac{1}{\tau^{(k)}}\equiv\frac{e}{h}\Big|\bm{E}_{\rm cn}\cdot\left(\left\langle 1^{(k)}|\hat{\bm{r}}|1^{(k)}\right\rangle-\left\langle 0^{(k)}|\hat{\bm{r}}|0^{(k)}\right\rangle\right)\Big|\,.$ (5)
Here, $\bm{E}_{\rm cn}$ is a characteristic value of the charge-noise induced electric field vector. If the field is spatially homogeneous throughout the QD region then, due to the symmetries of the confinement potential, the dominant contribution in $1/\tau$ is the one related to the $z$ component of $\bm{E}_{\rm cn}$. As in the case of the Rabi frequencies, attention should be paid to the trends and to the comparison between single- and three-hole spin qubits, rather than to the absolute values of the dephasing time scale, which depend on the assumed value of the charge-noise induced electric field.
The differences between the values of $\tau$ obtained in the single- and in the three-hole qubits are remarkable. Similarly to the case of the Rabi frequencies, the monotonic dependence of $\tau$ on the dot aspect ratio ($\omega_{x}$) that is obtained for the SHQ is replaced — in the case of the THQ — by a non-monotonic dependence, which varies qualitatively with the magnetic-field orientation (Fig. 4). In particular, we note that the dephasing time scale of the THQ is lower than that of the SHQ for a quasi-circular dot. However, the enhancement that the THQ allows in terms of Rabi frequency overcompensates such an effect, resulting in a higher value of a quality factor $Q_{x}\equiv f_{{\rm R},x}\tau$ (see below). As for the Rabi frequencies, the region corresponding to $\hbar\omega_{x}\approx 3.8\,$meV is characterized by strong variations of $\tau$, which we attribute to the effect on the ground state doublet of the narrow anticrossing affecting the two lowest excited doublets [76].